OpenStudy (anonymous):

Solve the quadratic equation by completing the square. –3x^2 + 9x = 1 A. 3/2±sqrt(6)/6 B. -3±sqrt(69)/3 C. 3±sqrt(-3)/3 D. 3/2±sqrt(69)/6

OpenStudy (helder_edwin):

first factor -3 from the left side

OpenStudy (helder_edwin):

then u have to add the coeficient of x halved and squared to both sides of the equation

OpenStudy (anonymous):

3x(-x + 3) = 1 I factors the -3 from the left side I don't get what you mean by the second part though and sorry for the late reply I was eating dinner.

OpenStudy (tkhunny):

–3x^2 + 9x = 1 I like the idea about the -3, but divide by it and forget about it, rather than factor it and have it complicate your life. x^2 - 3x = -1/3 NOW, we can worry about Completing the Square. What's next?

OpenStudy (anonymous):

x^2 = -1/3- 3x?

OpenStudy (tkhunny):

You seem to be solving by old rules. Your instructions are to "Complete the Square"

OpenStudy (anonymous):

x^2 = -1/3 - 1.5?

OpenStudy (anonymous):

It says to square it

OpenStudy (tkhunny):

It says, "Completing the Square". This is a known process. Not much room for making up what it is. You must exploit this relationship: \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (anonymous):

Hold on I think tk gave the basic idea to me.... i'm going to try it one sec

OpenStudy (solomonzelman):

 Hold